# Simulating Work Extraction in a Dinuclear Quantum Battery Using a Variational Quantum Algorithm

### v1

@@ I Introduction
These quantum thermal devices have been proposed using condensed matter physics, especially quantum battery, which is a device that stores energy in a quantum system to be used for some useful task. For instance, it was proposed using superconducting systems Strambini et al. (2020) and spin chains Le et al. (2018); Rossini et al. (2019); Zhao et al. (2021). Despite the advances in the theoretical framework, there are challenges to be overcome before we see a practical application of QBs, such as storing energy at room temperature, offering non-destructive access to the stored work, and a charge lifetime comparable with that of the conventional classical batteries.
@@
In this sense, low-dimensional metal complexes (LDMC) emerge as a promising platform for QBs implementations. These advanced materials exhibit unique properties, making them highly suitable for experimental quantum devices. For instance, their molecular structure shields them against environmental perturbations, e.g., high-temperature Reis et al. (2012); Čenčariková and Strečka (2020), magnetic fields Čenčariková and Strečka (2020); Cruz and Anka (2020); Souza et al. (2008), and pressure Cruz and Anka (2020); Cruz et al. (2017). Another useful feature is that LDMC presents an effective behavior of a two-qubit system that can hold stable quantum correlations above room temperature Cruz et al. (2016); Souza et al. (2009); Reis et al. (2012). In this context, some of us proposed quantum thermal devices that implement these advanced materials for quantum batteries at room temperature, showing that it is possible to extract work from genuine quantum correlations Cruz et al. (2022), and a mechanical model for the quantum Stirling cycle Cruz et al. (2023); Gomes de Paula et al. (2024).
@@
In this work, we explore a quantum battery model based on quantum discord using Variational Quantum Algorithms simulated using IBM Qiskit Team (2024). We simulate the optimal work extraction protocol proposed in Cruz et al. (2022) in both noiseless and noisy conditions and compare the numerical results with theoretical and experimental predictions. Our findings illustrate that, despite deviations induced by the noisy environments presented, the variational approach accurately captures the expected relationship between ergotropy and quantum correlations. Unlike the experimental method in Cruz et al. (2022), which infers stored energy via magnetic susceptibility measurements, our approach provides a computational framework for simulating and giving insights into optimizing work extraction in a dinuclear quantum battery model. Additionally, we expand the result presented in reference Cruz et al. (2022), analyzing how the noisy environment influences work fluctuations in the work extraction process, providing a perspective on the practical challenges of implementing quantum batteries with near-term quantum hardware. The results show that, although reference Cruz et al. (2022) demonstrated that the system is capable of energy storage at ambient conditions, our simulation of the work extraction protocol reveals satisfactory precision at lower temperatures, while the extraction accuracy at higher temperatures remains a limiting factor for its practical usability.
@@ II Theoretical foundations
In order to analyze the work extraction in a dinuclear quantum battery, it is important to establish the theoretical framework governing the working substance. In this context, this section aims to review the fundamental physical model of the system. First, the Hamiltonian framework of the dinuclear quantum battery will be presented, emphasizing the role of spin interactions and the contribution of the external field on its energy eigenstates. The discussion then advances to the concept of ergotropy, quantifying the amount of extractable work from the quantum system of interest. Finally, the optimal work extraction protocol is reviewed, expanding the consolidated references by adding an analysis of the associated energy fluctuations, providing the necessary groundwork for the subsequent quantum simulations. Thus, this section serves as the background for the variational quantum algorithm simulations, analyzing the performance of the system as a quantum battery under realistic conditions of a noisy quantum processor.
@@ II.1 Working Substance
Ref.Cruz et al. (2022) developed a theoretical model for a two-cell quantum battery based on a metal-organic copper (II) compound - Cu2(HCOO)4(HCOOH)2(C4H10N2). The reduced magnetic unit is formed by two metallic centers of Cu(II), with electronic configuration d9 and $s=1/2$, with a short intermolecular separation, characteristic of a carboxylate-based metal complex Cruz et al. (2016). Such property provides a nearly ideal realization of an isolated two-qubit system Yurishchev (2011); Gaita-Ariño et al. (2019); Moreno-Pineda et al. (2018), since their intramolecular interaction energy is extremely large when compared to their intermolecular ones Cruz et al. (2016); Yurishchev (2011); Gaita-Ariño et al. (2019); Moreno-Pineda et al. (2018), which shields the quantum properties of the system against high-temperature effects, enabling the system to store energy throughout quantum correlations at room temperature.
@@
$H=E_{0}(S_{1}^{z}+S_{2}^{z})+J(\vec{S}_{1}\cdot\vec{S}_{2}),$ (1)
where $S_{n}^{k}=(1/2)\sigma_{n}^{k}$, with $\sigma_{n}^{k}$ is the Pauli matrix of the i-th spin and $k\in\{x,y,z\}$. The first term is the Zeeman Hamiltonian, where $E_{0}=\mu_{B}g_{z}B_{z}$, $g_{z}$ is the isotropic Landé factor, $\mu_{B}$ is the Bohr magneton, and $B_{z}$ is the external magnetic field, describing the energy levels of each cell of the battery: $H_{0}\left|\epsilon_{i}\right\rangle=\epsilon_{i}\left|\epsilon_{i}\right\rangle$, with associated eigenvalues $\epsilon_{1}=-E_{0}$, $\epsilon_{2}=\epsilon_{3}=0$, and $\epsilon_{4}=E_{0}$, and respective eigenstates $\left|\epsilon_{1}\right\rangle=\left|00\right\rangle$, $\left|\epsilon_{2}\right\rangle=\left|10\right\rangle$, $\left|\epsilon_{3}\right\rangle=\left|01\right\rangle$, and $\left|\epsilon_{4}\right\rangle=\left|11\right\rangle$ defined by the local basis Cruz et al. (2022). The second term is the Heisenberg Hamiltonian, which accounts for the internal interaction between Cu(II) ions, while $J$ represents the magnetic coupling constant. The energy structure of this Hamiltonian is composed of a singlet state $\left|\beta_{-}\right\rangle$ with energy $E_{-}=-3J/4$ and a triplet-degenerate subspace $\{\left|\beta_{t}\right\rangle\}=\{\left|\beta_{+}\right\rangle,\left|00\right\rangle,\left|11\right\rangle\}$, with energy $E_{t}=J/4$ and $\left|\beta_{\pm}\right\rangle=(\left|01\right\rangle\pm\left|10\right\rangle)/2$.
@@ II.2 Ergotropy
The ergotropy can be written as the maximum work that can be extracted from a quantum system via unitary operation $V$:
$\begin{split}\mathcal{E}(T,B_{z})&=\mbox{max}_{U\in\mathcal{U}}\{\mathcal{W}(T,B_{z})\}\\ &=\operatorname{tr}[\rho(T,B_{z})H_{0}]-\mbox{min}_{U\in\mathcal{U}}\{\operatorname{tr}[U\rho(T,B_{z})U^{\dagger}H_{0}]\},\end{split}$ (2)
with the maximization/minimization is taken over the set $\mathcal{V}$ of all unitary operators. From this equation, it is clear that not all internal energy can be extracted as work, since $\mathcal{E}(T,B_{z})=E(\rho(T,B_{z}))-E_{0}$, where $E(\rho(T,B_{z}))$ is the internal energy of the initial system Allahverdyan et al. (2004); Niedenzu et al. (2019).
It is worth noting that we are assuming that the initial state of the system is a thermal state given by $\rho(T,B_{z})=e^{-H/k_{B}T}/tr[e^{-H/k_{B}T}]$, which is diagonal in the coupled basis $\{\left|\varrho_{i}\right\rangle\}$. Thus, it is a complete passive state, which no energy can be extracted from. However, we consider the density matrix in the uncoupled basis of the self-Hamiltonian of the battery $\{\left|\epsilon_{j}\right\rangle\}$, from which it takes the form of an active X-state, in which a finite amount of energy can be extracted:
@@ II.3 Optimal Work Extraction
In this scenario, Ref.Cruz et al. (2022) introduces the evaluation of the ergotropy through an experimental measure of a macroscopic property, the magnetic susceptibility $\chi(T)$. This is achievable for the regime of magnetic susceptibility, $E_{0}<<k_{B}T$. In this case, we can write the ergotropy as a function of the temperature as
@@
in terms of $\chi(T)$ given by the Bleaney-Bowers equation s Bleaney and Bowers (1952)
@@
where $N_{A}$ is the Avogrado number. In this regard, Eq. (4) can be written directly proportional to the quantum discord of the system:
@@
is the Schatten one-norm quantum discord Cruz et al. (2016).
The work extraction protocol is characterized by the unitary operator $U=\sum_{i}\left|\varepsilon_{i}\right\rangle\left\langle\varrho_{i}\right|$ Allahverdyan et al. (2004), such that the initial active state becomes a passive state, $\rho(T,B_{z})\rightarrow\sigma(T,B_{z})=U\rho(T,B_{z})U^{\dagger}$, where $[\sigma(T,B_{z}),H_{0}]=0$. In the studied case, in the low-temperature limit $k_{B}T\ll J$, the maximum ergotropy state is given by $\rho=\left|\beta_{-}\right\rangle\left\langle\beta_{-}\right|$, while the empty state, with zero ergotropy, is given by $\sigma=\left|00\right\rangle\left\langle 00\right|$. This optimal work extraction is experimentally achievable in the thermal range of $T<83$ K Cruz et al. (2022). Therefore, in the quantum computing framework, the unitary operator responsible for this operation can be implemented by a sequence of logical gates as follows:
@@
Thus this optimal work extraction can be tested by exploring a quantum computing simulation as a proof of principle for the proposed quantum battery model using a quantum computer to implement this work extraction protocol.
@@ II.4 Protocol Precision
Before presenting the quantum computing framework for the energy extraction protocol, we briefly expand the result presented in reference Cruz et al. (2022) by discussing a figure of merit for analyzing the charging/discharging process in quantum batteries. In Ref.Friis and Huber (2018), the authors proposed a way to evaluate the precision in the charging process by analyzing the average energy increase after the unitary operation. To charge an initial thermal state $\tau$ by unitary operation, such that $\rho=U\tau U^{\dagger}$, an amount of average energy must be increased, leading to $\Delta E=E(\rho)-E(\tau)$, where $E(\rho)=\operatorname{tr}[\rho H_{0}]$. The precision of this process is evaluated by an increase in the standard deviation of the system Hamiltonian:
@@
where $V(\rho)=(\Delta H(\rho))^{2}=\operatorname{tr}(H_{0}^{2}\rho)-(\operatorname{tr}(H_{0}\rho))^{2}$. The smaller variance corresponds to a higher precision of the charging/discharging process. Thus, Eq. (9) serves as a parameter for estimating the confidence in the implemented energy extraction protocol Friis and Huber (2018).
@@ III Energy extraction protocol using a variational quantum algorithm
Simulating quantum properties in complex systems, such as metal complexes, is a significant challenge in computational chemistry and condensed matter physics Liu and Cui (2018). These systems exhibit intricate electronic interactions that often demand methods capable of capturing their quantum nature beyond classical computational capabilities Zobel and González (2021). Variational Quantum Algorithms (VQAs) Cerezo et al. (2021); McClean et al. (2016) have emerged as a promising approach to address this complexity, leveraging the potential of quantum computers to simulate quantum systems more efficiently Kandala et al. (2017), especially considering the limitations of NISQ-era computers Preskill (2018).
VQAs operate within a hybrid computational framework, combining the strengths of classical and quantum devices. The key idea is to encode the quantum properties of the problem into a parametrized quantum circuit (ansatz) Cerezo et al. (2021), representing the state of the system under study. This ansatz is iteratively optimized by minimizing a cost function that encodes the system’s physical properties, such as its energy. Classical optimization algorithms guide this iterative process, adjusting the parameters to find the extrema of the cost function, which corresponds to the problem’s solution. Recent studies have introduced variational quantum algorithms to study work extraction in quantum systems Medina et al. (2024); Hoang et al. (2024).
Conversely, in the context of dinuclear metal complexes, VQAs can be used to investigate the thermal properties derived from the Hamiltonian of the system das Neves Silva et al. (2024); Verdon et al. (2019). This involves the preparation of thermal states that encode information about the properties of the complex at finite temperatures. Simulating these thermal states makes it possible to extract quantities such as specific heat and magnetic susceptibility das Neves Silva et al. (2024).

### v2

@@ I Introduction
These quantum thermal devices have been proposed using condensed matter physics, especially the quantum battery, which is a device that stores energy in a quantum system to be used for some useful task. For instance, it was proposed using superconducting systems Strambini et al. (2020) and spin chains Le et al. (2018); Rossini et al. (2019); Zhao et al. (2021). Despite the advances in the theoretical framework, there are challenges to be overcome before we see a practical application of QBs, such as storing energy at room temperature, offering non-destructive access to the stored work, and a charge lifetime comparable with conventional classical batteries.
@@
Recent studies have demonstrated the experimental viability of QBs using distinct physical platforms, such as microcavities enclosing molecular dyes Quach et al. (2022), star-topology NMR spin systems Joshi and Mahesh (2022), and multi-layered organic microcavities in scalable room-temperature implementation Hymas et al. (2025). In parallel, low-dimensional metal complexes (LDMC) emerge as a promising platform for QBs implementations. These advanced materials exhibit unique properties, making them highly suitable for experimental quantum devices. For instance, their molecular structure shields them against environmental perturbations, e.g., high-temperature Reis et al. (2012); Čenčariková and Strečka (2020), magnetic fields Čenčariková and Strečka (2020); Cruz and Anka (2020); Souza et al. (2008), and pressure Cruz and Anka (2020); Cruz et al. (2017). Another useful feature is that LDMC presents an effective behavior of a two-qubit system that can hold stable quantum correlations above room temperature Cruz et al. (2016); Souza et al. (2009); Reis et al. (2012). In this context, some of us proposed quantum thermal devices that implement these advanced materials for quantum batteries at room temperature, showing that it is possible to extract work from genuine quantum correlations Cruz et al. (2022), and a mechanical model for the quantum Stirling cycle Cruz et al. (2023); Gomes de Paula et al. (2024). The core models used in these works Cruz et al. (2022, 2023, 2016); Souza et al. (2009); Gomes de Paula et al. (2024); Reis et al. (2012) are based on dinuclear metal complexes, i.e., a simple magnetic structure composed of two coupled metallic centers, composed by magnetic ions linked directly or through intermediate chemical groups (ligands) Reis (2013).
In this scenario, theoretical investigations become increasingly crucial to guide and complement these experimental progresses Razzoli et al. (2025). For this purpose, quantum simulations offer a powerful tool able to explore the dynamical behavior of these systems under realistic conditions, allowing one to assess the impact of decoherence, dissipation, and finite-temperature effects on device performance Georgescu et al. (2014); Lloyd (1996). In this context, Variational Quantum Algorithms (VQAs) Cerezo et al. (2021); McClean et al. (2016) have emerged as a promising approach to address this complexity, leveraging the potential of quantum computers to simulate quantum systems more efficiently Kandala et al. (2017), especially considering the limitations of NISQ-era devices Preskill (2018). Moreover, the potentialities of quantum variational schemes were explored in ref. Razzoli et al. (2025), where the authors simulated a cyclic solid-state quantum battery on superconducting IBM quantum machines.
@@
In this work, we explore a quantum battery model based on quantum discord using Variational Quantum Algorithms simulated using IBM Qiskit Team (2024). We simulate the optimal work extraction protocol proposed in Cruz et al. (2022), for a two-cell (dinuclear) quantum battery, in both noiseless and noisy conditions, and compare the numerical results with theoretical and experimental predictions. Our findings illustrated that, despite deviations induced by the noisy environments presented, the variational approach accurately captures the expected relationship between ergotropy and quantum correlations. Unlike the experimental method in Cruz et al. (2022), which infers stored energy via magnetic susceptibility measurements, our approach provides a computational framework for simulating and giving insights into optimizing work extraction in a dinuclear quantum battery model. Additionally, we expand the result presented in reference Cruz et al. (2022), analyzing how the noisy environment influences work fluctuations in the work extraction process, providing a perspective on the practical challenges of implementing quantum batteries with near-term quantum hardware. The results show that, although reference Cruz et al. (2022) demonstrated that the system is capable of energy storage at ambient conditions, our simulation of the work extraction protocol reveals satisfactory precision at lower temperatures, while the extraction accuracy at higher temperatures remains a limiting factor for its practical usability.
@@ II Theoretical foundations
In order to analyze the work extraction in a dinuclear quantum battery, it is important to establish the theoretical framework governing the working substance. In this context, this section aims to review the fundamental physical model of the system. First, the Hamiltonian framework of the dinuclear quantum battery will be presented, emphasizing the role of spin interactions and the contribution of the external field to its energy eigenstates. The discussion then advances to the concept of ergotropy, quantifying the amount of extractable work from the quantum system of interest. Finally, the optimal work extraction protocol is reviewed, expanding the consolidated references by adding an analysis of the associated energy fluctuations, providing the necessary groundwork for the subsequent quantum simulations. Thus, this section serves as the background for the variational quantum algorithm simulations, analyzing the performance of the system as a quantum battery under realistic conditions of a noisy quantum processor.
@@ II.1 Working Substance
Reference Cruz et al. (2022) developed a theoretical model for a two-cell quantum battery based on a metal-organic copper (II) compound - Cu2(HCOO)4(HCOOH)2(C4H10N2). The reduced magnetic unit is formed by two metallic centers of Cu(II), with electronic configuration d9 and $s=1/2$, with a short intermolecular separation, characteristic of a carboxylate-based metal complex Cruz et al. (2016). Such property provides a nearly ideal realization of an isolated two-qubit system Yurishchev (2011); Gaita-Ariño et al. (2019); Moreno-Pineda et al. (2018), since their intramolecular interaction energy is extremely large when compared to their intermolecular ones Cruz et al. (2016); Yurishchev (2011); Gaita-Ariño et al. (2019); Moreno-Pineda et al. (2018), which shields the quantum properties of the system against high-temperature effects, enabling the system to store energy throughout quantum correlations at room temperature.
@@
$\mathcal{H}=E_{0}(S_{1}^{z}+S_{2}^{z})+J(\vec{S}_{1}\cdot\vec{S}_{2}),$ (1)
where $S_{n}^{k}=(1/2)\sigma_{n}^{k}$, with $\sigma_{n}^{k}$ is the Pauli matrix of the $n-th$ spin and $k\in\{x,y,z\}$. The first term is the Zeeman Hamiltonian $\mathcal{H}_{0}=E_{0}(S_{1}^{z}+S_{2}^{z})$, where $E_{0}=\mu_{B}g_{z}B_{z}$, $g_{z}$ is the isotropic Landé factor, $\mu_{B}$ is the Bohr magneton, and $B_{z}$ is the external magnetic field, describing the energy levels of each cell of the battery: $\mathcal{H}_{0}\left|\epsilon_{i}\right\rangle=\epsilon_{i}\left|\epsilon_{i}\right\rangle$, with associated eigenvalues $\epsilon_{1}=-E_{0}$, $\epsilon_{2}=\epsilon_{3}=0$, and $\epsilon_{4}=E_{0}$, and respective eigenstates $\left|\epsilon_{1}\right\rangle=\left|00\right\rangle$, $\left|\epsilon_{2}\right\rangle=\left|10\right\rangle$, $\left|\epsilon_{3}\right\rangle=\left|01\right\rangle$, and $\left|\epsilon_{4}\right\rangle=\left|11\right\rangle$ defined by the local basis Cruz et al. (2022). The second term is the Heisenberg Hamiltonian, which accounts for the internal interaction between Cu(II) ions, while $J$ represents the magnetic coupling constant. The energy structure of this Hamiltonian is composed of a singlet state $\left|\beta_{-}\right\rangle$ with energy $E_{-}=-3J/4$ and a triplet-degenerate subspace $\{\left|\beta_{t}\right\rangle\}=\{\left|\beta_{+}\right\rangle,\left|00\right\rangle,\left|11\right\rangle\}$, with energy $E_{t}=J/4$ and $\left|\beta_{\pm}\right\rangle=(\left|01\right\rangle\pm\left|10\right\rangle)/\sqrt{2}$.
@@ II.2 Ergotropy
The ergotropy can be written as the maximum work that can be extracted from a quantum system via unitary operation $U$:
$\begin{split}\mathcal{E}(T,B_{z})&=\operatorname{tr}[\rho(T,B_{z})\mathcal{H}_{0}]-\mbox{min}_{U\in\mathcal{U}}\{\operatorname{tr}[U\rho(T,B_{z})U^{\dagger}\mathcal{H}_{0}]\},\end{split}$ (2)
where the minimization is taken over the set $\mathcal{U}$ of all unitary operators. From this equation, it is clear that not all internal energy ($E$) can be extracted as work, since $\mathcal{E}(T,B_{z})=E(\rho(T,B_{z}))-E_{0}$, where $E(\rho(T,B_{z}))$ is the internal energy of the initial system Allahverdyan et al. (2004); Niedenzu et al. (2019). It was shown that, by ordering the eigenvalues of the self-Hamiltonian in increasing order, $\epsilon_{1}\leq\epsilon_{2}\leq\cdots\leq\epsilon_{N}$, and the eigenvalues of the density matrix of the system in decreasing order, $\varrho_{1}\geq\varrho_{1}\geq\cdots\geq\varrho_{1}$, we can write the ergotropy as Allahverdyan et al. (2004)
$\mathcal{E}(T,B_{z})=\sum_{i,j}^{N,N}\varrho_{i}\epsilon_{j}(|\langle\varrho_{i}|\epsilon_{j}\rangle|^{2}-\delta_{i,j}).$ (3)
It is worth noting that we are assuming that the initial state of the system is a thermal state given by $\rho(T,B_{z})=e^{-\mathcal{H}/k_{B}T}/tr[e^{-\mathcal{H}/k_{B}T}]$, which is diagonal in the coupled basis $\{\left|\varrho_{i}\right\rangle\}$. Thus, it is a complete passive state, from which no energy can be extracted. However, we consider the density matrix in the uncoupled basis of the self-Hamiltonian of the battery $\{\left|\epsilon_{j}\right\rangle\}$, from which it takes the form of an active X-state, in which a finite amount of energy can be extracted:
@@ II.3 Optimal Work Extraction
In this scenario, Ref.Cruz et al. (2022) introduces the evaluation of the ergotropy through an experimental measure of a macroscopic property, the magnetic susceptibility $\chi(T)$. This is achievable for the regime of magnetic susceptibility, $E_{0}<<k_{B}T$. In this limit, using Eq.(3), we can write the ergotropy as a function of the temperature as Cruz et al. (2022)
@@
in terms of $\chi(T)$ given by the Bleaney-Bowers equation  Bleaney and Bowers (1952)
@@
where $N_{A}$ is the Avogadro number. In this regard, Eq. (5) can be written directly proportional to the quantum discord of the system Cruz et al. (2022):
@@
is the Schatten one-norm quantum discord Cruz et al. (2016)  [^1] More details on the general form of the ergotropy and quantum discord, and the derivation of Eqs.(5-8) can be found in the appendix of Ref.Cruz et al. (2022)..
The work extraction protocol is characterized by the unitary operator $U=\sum_{i}\left|\varepsilon_{i}\right\rangle\left\langle\varrho_{i}\right|$ Allahverdyan et al. (2004), such that the initial active state becomes a passive state, $\rho(T,B_{z})\rightarrow\sigma(T,B_{z})=U\rho(T,B_{z})U^{\dagger}$, where $[\sigma(T,B_{z}),\mathcal{H}_{0}]=0$. In the studied case, in the low-temperature limit $k_{B}T\ll J$, the maximum ergotropy state is given by $\rho=\left|\beta_{-}\right\rangle\left\langle\beta_{-}\right|$, while the empty state, with zero ergotropy, is given by $\sigma=\left|00\right\rangle\left\langle 00\right|$. This optimal work extraction is experimentally achievable in the thermal range of $T<83$ K Cruz et al. (2022). At this temperature, the system reaches the pure state regime. Thus, the $83$ K temperature represents a thermodynamic threshold which marks the transition from a regime of maximized quantum resources Cruz et al. (2016). Above this temperature, the thermal energy becomes significant to populate excited states, causing the loss of the purely ground-state occupancy. The emergence of this threshold temperature is attributed to the magnetic coupling characteristics inherent to dinuclear metal complexes, particularly those based on carboxylate groups Cruz et al. (2016, 2022). In such materials, a strong metal-to-metal magnetic interaction, mediated through carboxylate bridges with a syn–syn conformation Cruz et al. (2016), leads to a significant energy gap between the singlet ground state and the first excited triplet states Cruz et al. (2022). Therefore, below this temperature, the system predominantly occupies the maximally entangled singlet ground state, maximizing quantum correlations and leading to an optimal work extraction in a quantum battery model Cruz et al. (2022).
Therefore, in the quantum computing framework, the unitary operator responsible for this operation can be implemented by a sequence of logical gates as follows:
@@
This sequence first applies a CNOT gate with qubit 1 as control and qubit 2 as target ($CNOT_{12}$), followed by a Hadamard gate on qubit 1 ($H_{1}$). Thus, this optimal work extraction can be tested by exploring a quantum computing simulation as a proof of principle for the proposed quantum battery model using a quantum computer to implement this work extraction protocol.
@@ II.4 Protocol Precision
Before presenting the quantum computing framework for the energy extraction protocol, we briefly expand the result presented in reference Cruz et al. (2022) by discussing a figure of merit for analyzing the charging/discharging process in quantum batteries. In Ref.Friis and Huber (2018), the authors proposed a way to evaluate the precision in the charging process by analyzing the average energy increase after the unitary operation. To charge an initial thermal state $\tau$ by unitary operation, such that $\rho=U\tau U^{\dagger}$, an amount of average energy must be increased, leading to $\Delta E=E(\rho)-E(\tau)$, where $E(\rho)=\operatorname{tr}[\rho\mathcal{H}_{0}]$. The precision of this process is evaluated by an increase in the standard deviation of the system Hamiltonian:
@@
where $V(\rho)=(\Delta\mathcal{H}_{0}(\rho))^{2}=\operatorname{tr}(\mathcal{H}_{0}^{2}\rho)-(\operatorname{tr}(\mathcal{H}_{0}\rho))^{2}$. The smaller variance corresponds to a higher precision of the charging/discharging process. Thus, Eq. (10) serves as a parameter for estimating the confidence in the implemented energy extraction protocol Friis and Huber (2018).
@@ III Energy extraction protocol using a variational quantum algorithm
Simulating quantum properties in complex systems, such as metal complexes, is a significant challenge in computational chemistry and condensed matter physics Liu and Cui (2018). These systems exhibit intricate electronic interactions that often demand methods capable of capturing their quantum nature beyond classical computational capabilities Zobel and González (2021). An alternative with the potential to avoid this problem is the Variational Quantum Algorithms (VQAs): a hybrid computational framework that combines the strengths of classical and quantum devices.
Variational Quantum Algorithms (VQAs) represent one of the most promising approaches to harness the potential of intermediate-scale quantum devices, known as NISQ (Noisy Intermediate-Scale Quantum) devices Bharti et al. (2022). In this context, VQAs have been widely applied to solve a variety of relevant problems, such as computing energy spectra in molecular systems Kandala et al. (2017), combinatorial optimization Farhi et al. (2014), and training machine learning models Mitarai et al. (2018). Among these applications, quantum system simulation stands out as a field of great interest, encompassing everything from modeling interactions in strongly correlated systems McArdle et al. (2020); Kandala et al. (2017) to exploring dynamics in out-of-equilibrium systems Yuan et al. (2019).
In a typical VQA, the main idea is to encode the quantum properties of the problem into a parameterized quantum circuit (ansatz) Cerezo et al. (2021), which represents the state of the system under study. This ansatz is iteratively optimized by minimizing a cost function that encodes the system’s physical properties, such as its energy. Classical optimization algorithms guide this iterative process, adjusting the parameters to find the extrema of the cost function, which correspond to the solution of the problem.
A well-known example of this approach is the simulation of the ground state energy of a simple molecule like $H_{2}$. In this case, a quantum circuit prepares a trial state, the system’s energy is measured, and a classical optimizer adjusts the circuit parameters to minimize that energy. Repetition of this process converges the algorithm toward the ground state of the molecule, demonstrating the effectiveness of VQAs in capturing quantum behavior even on noisy quantum devices with few qubits. This simple case has already been demonstrated experimentally Peruzzo et al. (2014).
In a previous work, we used a VQA, the Variational Quantum Thermalizer algorithm (VQT), to investigate the thermal properties of dinuclear metal complexes based on the Hamiltonian of the system das Neves Silva et al. (2024). This involves the preparation of thermal states that encode information about the properties of the complex at finite temperatures. Simulating these thermal states makes it possible to extract quantities such as specific heat and magnetic susceptibility.
Several alternative methods for preparing thermal states exist, including quantum Metropolis sampling Temme et al. (2011), Grover-based algorithms Chiang and Wocjan (2010), and Thermal Pure Quantum (TPQ) states Sugiura and Shimizu (2013). However, they often require deep circuits or large systems, making them unsuitable for NISQ devices. Variational methods are more compatible with near-term hardware. Among them, VQT stands out by optimizing a physically motivated cost function, the quantum free energy, which naturally balances energy minimization and entropy maximization. This avoids simulating non-unitary dynamics (as in variational imaginary-time evolution McArdle et al. (2019)) and enables direct convergence to thermal equilibrium.